# DySLIM: Dynamics Stable Learning by Invariant Measure for Chaotic Systems

### v1

@@ 4 Main Idea and Methods
To tackle the issue of distribution shift, we propose to focus on systems’ invariant measure preservation. Specifically, many systems of interest have some measure $\mu^{*}$ supported on an attractor that is invariant to the transformation ${\mathcal{S}}$ (Tucker, 2002; Luzzatto et al., 2005; Ferrario, 2008; Weinan & Liu, 2002; Hawkins, 2021).
We cast our learning problem as finding parameters $\theta$ such that a surrogate ${\mathcal{S}}_{\theta}$ preserves $\mu^{*}$ while approximating ${\mathcal{S}}$ locally, which defines the following constrained optimization:
@@
where ${\mathcal{L}}(\theta)$ is the short-term loss, and $\mu_{\theta}^{*}$ is the invariant measure of ${\mathcal{S}}_{\theta}$, i.e., $({\mathcal{S}}_{\theta})_{\sharp}\mu_{\theta}^{*}=\mu_{\theta}^{*}$, which we assume exists [^5] This is a key hypothesis in our methodology. Otherwise this assumption can be enforced in ${\mathcal{S}}_{\theta}$ by adding a potential term, as done in Li et al. (2022).. Solving this constrained optimization inherently alleviates the distribution shift problem. Since trivial solutions exist for measure preservation, e.g., if ${\mathcal{S}}_{\theta}$ is the identity, the trajectory matching component ${\mathcal{L}}(\theta)$ of this constrained objective is necessary for producing useful surrogates.
@@
Our choice of measure distance needs to satisfy several desiderata, namely it should: i) respect the underlying geometry of ${\mathcal{U}}$ and support comparison between measures with non-overlapping supports, ii) admit an unbiased, sampled-based estimator, iii) have low computational complexity with respect to the system dimension and number of samples, iv) entail convergence properties on the space of measures defined on ${\mathcal{U}}$ (informally, $\textsc{D}(\mu_{\theta}^{*},\mu^{*})\rightarrow 0\implies\mu_{\theta}^{*}\rightarrow\mu^{*}$), and v) enjoy parametric rates of estimation (i.e., sampling error $|\textsc{D}-\widehat{\textsc{D}}|$ is independent of system dimension).
@@
Some popular notions of distance / divergence from statistical learning theory include the Kullback-Leibler and Hellinger. However, these do not take into account the distance metric of the space on which the distributions are defined (Feydy et al., 2019; Genevay et al., 2018) and in some cases are undefined for non-overlapping supports.
@@
In contrast, Integral Probability Metrics (IPMs; Müller (1997)) represent a general purpose tool for comparing for two samples. Among the class of IPMs, the Maximum Mean Discrepancy (MMD; Gretton et al. (2012)) stands out as it has a closed form expression and satisfies all our requirements described above. Deferring several details about the MMD to Appendix B, we define it here as,
$\mathrm{MMD}^{2}(\mu^{*},\mu_{\theta}^{*})=\mathbb{E}_{{\bm{u}},{\bm{u}}^{\prime}\sim\mu^{*}}[\kappa({\bm{u}},{\bm{u}}^{\prime})]+\mathbb{E}_{{\bm{v}},{\bm{v}}^{\prime}\sim\mu_{\theta}^{*}}[\kappa({\bm{v}},{\bm{v}}^{\prime})]-2\mathbb{E}_{{\bm{u}}\sim\mu^{*},{\bm{v}}\sim\mu_{\theta}^{*}}[\kappa({\bm{u}},{\bm{v}})],$ (10)
@@
where $\kappa:{\mathcal{U}}\times{\mathcal{U}}\rightarrow\mathbb{R}$ is a kernel. [^6] In practice, the choice of kernel has important practical implications (Liu et al., 2020; Schrab et al., 2023), and many kernels $\kappa_{\sigma}$ are controlled by a bandwidth hyperparameter $\sigma$ that should be tuned. For two sets of $n$ samples $\{{\bm{u}}^{(i)}\}_{i=1}^{n}\sim\mu^{*}$ and $\{{\bm{v}}^{(i)}\}_{i=1}^{n}\sim\mu_{\theta}^{*}$, Equation 10 admits an unbiased estimator (Gretton et al., 2012):
$\widehat{\mathrm{MMD}}^{2}(\mu^{*},\mu_{\theta}^{*})=\frac{1}{n(n-1)}\Bigg{[}\sum_{i=1}^{n}\sum_{j\neq i}^{n}[\kappa({\bm{u}}^{(i)},{\bm{u}}^{(j)})]+\sum_{i=1}^{n}\sum_{j\neq i}^{n}[\kappa({\bm{v}}^{(i)},{\bm{v}}^{(j)})]\Bigg{]}-\frac{2}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}[\kappa({\bm{u}}^{(i)},{\bm{v}}^{(j)})],$
@@
This estimator can be easily computed in $O(n^{2})$ operations. Additionally, the MMD entails convergence properties on the space of measures defined on ${\mathcal{U}}$. That is, it metrizes weak convergence (Simon-Gabriel et al., 2023), and, for characteristic kernels, we have $\mathrm{MMD}(\mu^{*},\mu_{\theta}^{*})=0\iff\mu^{*}=\mu_{\theta}^{*}$ (Sriperumbudur et al., 2010). The MMD also enjoys parametric rates of estimation, with $O(1/\sqrt{n})$ sampling error (Gretton et al., 2006; Tolstikhin et al., 2016).
@@
Using the approximation in Equation 11 in the context of a stochastic optimization pipeline requires that we estimate this term with a potentially small batch-size. In fact, besides some simple systems, one typically can only afford small batch sizes, which means we may not be fully capturing both $\mu^{*}$ and $\mu_{\theta}^{*}$. Although our choice of regularization loss comes with an unbiased estimator with parametric rates of error, in small regimes, estimation error can still diminish its effectiveness at providing a meaningful signal.
@@
$\mathrm{MMD}^{2}(({\mathcal{S}}^{k})_{\sharp}\mu^{*},({\mathcal{S}}_{\theta}^{k})_{\sharp}\mu^{*})=\mathbb{E}_{{\bm{u}},{\bm{v}}\sim\mu^{*}}[\kappa({\mathcal{S}}^{k}({\bm{u}}),{\mathcal{S}}^{k}({\bm{v}}))+\kappa({\mathcal{S}}_{\theta}^{k}({\bm{u}}),{\mathcal{S}}_{\theta}^{k}({\bm{v}}^{\prime}))-2\kappa({\mathcal{S}}^{k}({\bm{u}}),{\mathcal{S}}_{\theta}^{k}({\bm{v}})],$ (12)
which can be estimated by extracting a subset of initial conditions, time-evolving them $k$ steps, and then computing the estimator on the time-evolved samples. We point out that this is equivalent to conditioning the loss on the initial conditions, hence the name conditional regularization.
@@
^$\widehat{{\mathcal{L}}}_{\lambda}^{\textsc{D}}(\theta)=\widehat{{\mathcal{L}}}^{\text{obj}}(\theta)+\lambda_{1}\widehat{\textsc{D}}(\mu^{*},({\mathcal{S}}_{\theta}^{\ell})_{\sharp}\mu^{*})+\lambda_{2}\widehat{\textsc{D}}(({\mathcal{S}}^{\ell})_{\sharp}\mu^{*},({\mathcal{S}}_{\theta}^{\ell})_{\sharp}\mu^{*}),$ (13)
@@
For the measure distance $\widehat{\textsc{D}}$ in Equation 13, we use $\widehat{\mathrm{MMD}}^{2}$ and define $\kappa_{\bm{\sigma}}$ as a mixture of rational quadratic kernels (Rasmussen et al., 2006) with varying bandwidths (Li et al., 2015):
@@ 5 Experiments
Evaluation
We evaluate models both for their ‘short-term’ predictive ability and ‘long-run’ stability. The former is measured by a cosine similarity statistic between true and predicted trajectories. The latter is measured by system-specific metrics (see Wan et al. (2023a)) capturing the distributional similarity between true and generated trajectories, along with their visual inspection . In particular, we use the Sinkhorn Divergence (SD) (Genevay et al., 2018) between the empirical distributions of ground truth and predicted trajectories at various time steps to quantify distributional overlap. We also use the mean energy log ratio (MELR), which measures the average deviation of the energy at each Fourier mode of the generated snapshots when compared to the ground truth. We also consider its weighted variant (MELRw), which up-weights modes with higher energy, in particular the low-frequency modes. We also use the mean of the Frobenius norm of covariance matrix (covRMSE) that measures the spatial statistical properties of the generated samples. We additionally compute point-wise Wasserstein metrics. Finally, we consider a time correlation metric (TCM) which provides a measure of temporal behavior, in contrast to most of the metrics above which are snapshot-based. For more detail on and precise definitions of the evaluation criteria, see Appendix C.
@@ 5.1 Lorenz 63
Figure 1 (b) and (c) demonstrates the improved stability from adding invariant measure regularization to the different training objectives considered in this paper. In particular, Figure 1 (b) demonstrates improved long-term statistics, as the distribution of points for models trained with the regularized objectives are closer to that of ground truth trajectories compared to those from models trained with unregularized objectives. In addition, Figure 1 (c) shows that by incorporating the regularization we obtain the added benefits of improving the short-term model prediction, with longer de-correlation times. For further results with different metrics, see Appendix F.1.
@@ 5.2 Kuramoto-Sivashinsky
As with the Lorenz 63 experiments, from Figure 2 we observe better short-term predictions and improved long-term stability as measured by lower SD between ground truth and predicted trajectory distributions. In Figure 1 (a), we see example trajectories that highlight the difference between models trained with and without regularized objectives. For the Curr objective, models often diverge and produce numerical instability, while for the Pfwd objective, models deviate from the attractor. In contrast, the regularized versions of these objectives yield more stable models that remain on the correct attractor manifold.
@@ 5.3 Kolmogorov Flow
Finally, we study chaotic 2D fluid flow defined by the Navier-Stokes equations with Kolmogorov forcing. As above, information about the PDE and experiment setups are detailed in Appendix E.3. For the Navier-Stokes example the SD becomes non-discriminative due to the high-dimension of the state space, so we rely on the other metrics outlined above.
Figure 3 (left) shows a typical behavior of the unrolled trained models for a given initial condition: the baselines become highly dissipative and quickly veer towards the mean, whereas DySLIM greatly improves the long-term behavior. Figure 3 (right) shows a similar results to those in the other experiments: the short-term behavior of the solution is enhanced by the regularization (see Figure 8 for further comparisons). We point out that the curriculum training is often worse due to more stringent memory requirement that prevent us from unrolling the model during training for longer time-horizons.
As ablation studies, we sweep over different batch sizes and learning rates (see Appendix E.3 for the specific details of the experiment’s setup.) The results are summarized in Table 1, which shows that models trained with DySLIM either have an edge or remain competitive across the spectrum of different learning rates and batch sizes considered. However, as batch size and learning rate increase, the behavior of the model trained with DySLMI remains consistent, whereas the models trained only using the original objective deteriorate quickly.
@@
Figure 3: (left) Sample reference and predicted trajectory across models trained on the Kolmogorov Flow data using the Curr and Pfwd objectives, together with the regularized versions. (right) Evolution of the cosine similarity with time for Curr and Pfwd objectives with and without regularization. The solid line is the median among $160$ runs ($32$ trajectories for each of the $5$ random seeds), and the shaded regions corresponds to the second and third quartile. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = 128 and learning rate = 5e-4).
Table 1: Kolmogorov flow: metrics for pushforward, curriculum, and 1-step objectives without and with regularization ($\lambda_{1}=0$, $\lambda_{2}=100$). Boldface numbers indicate that the metric is improved by our regularization. See Table 2 for ablation studies on varying batch sizes and learning rates.
Baseline Batch size Learning rate
MELR↓
(×$\times 10^{-2}$)
MELRw↓
(×$\times 10^{-2}$)
covRMSE↓
(×$\times 10^{-2}$)
Wass1↓
(×$\times 10^{-2}$)
TCM↓
(×$\times 10^{-2}$)
@@ 6 Related work
Classical reduced order model (ROM) methods build surrogates by identifying low-dimensional linear approximation spaces tailored to representing target system states. Such spaces are usually derived from data samples (Aubry et al., 1988; Chinesta et al., 2011; Barrault et al., 2004; Amsallem et al., 2012), and ROMs are obtained by projecting the system equations onto the approximation space (Galerkin, 1915). Although these methods inherently leverage the linear behavior of underlying dynamics, they have been recently extended to handle mildly non-linear dynamics (Willcox, 2006; Astrid et al., 2008; Chaturantabut & Sorensen, 2010; Geelen et al., 2022; Ayed et al., 2019). However, their performance deteriorate rapidly for highly non-linear advection-dominated systems, such as KS and Kolmogorov flow (Peherstorfer, 2022)
Hybrid Physics-ML
@@
More recent methods hybridize classical numerical methods with contemporary data-driven deep learning techniques (Mishra, 2018; Bar-Sinai et al., 2019; Kochkov et al., 2021; List et al., 2022; Bruno et al., 2021; Frezat et al., 2022; Dresdner et al., 2022; Boral et al., 2023). These approaches learn corrections to numerical schemes from high-resolution simulation data, resulting in fast, low-resolution methods with high-accuracy. However, they require knowledge of the underlying PDE.
Pure ML-surrogates and Stabilization Techniques
@@
Recent works have focused on training on short-term trajectories while using recurrent networks (Vlachas et al., 2018) and reservoir computing techniques (Vlachas et al., 2020). Other approaches seek to regularize the training stage by leveraging properties of the systems. Such stabilization can be achieved by incorporating noise (Sanchez-Gonzalez et al., 2020), which can be induced by the learned model (Brandstetter et al., 2022); by back-propagating the gradient along many time steps (Um et al., 2020), and learning the dynamics on a latent space (Stachenfeld et al., 2022; Serrano et al., 2023), while promoting smoothness in the latent space (Wan et al., 2023b). Or by using a generative teacher network (Lamb et al., 2016), or leveraging an approximate inertial form (Lu et al., 2017). Finally, a somewhat related stabilization method, introduced in Wang et al. (2014); Blonigan et al. (2018), develops a shadowing technique for sensitivity analysis of long-term averaged gradients.
Operator Learning
@@
Neural operators seek to learn the integro-differential operators directly, without explicit PDE-informed components.These methods often leverage classical fast-methods (Li et al., 2021; Tran et al., 2021; Fan et al., 2019; Li et al., 2020), or approximation-theoretic structures (Lu et al., 2021) to achieve computational efficiency. Some of these techniques have been extended to handle dissipative systems (Li et al., 2022) by hard-coding a dissipative term at both training and inference time.
Learning Invariant Measures
@@
Botvinick-Greenhouse et al. (2023) use an Eulerian approach to learn dynamics on invariant measures for low-dimension ODEs using the Feynman-Kac formula coupled with PDE-constrained optimization. In concurrent work, Jiang et al. (2023) use neural operators with measure matching of handcrafted statistics, using optimal transport. However, the approach was only applied to small systems with low-dimensional attractors, and it is not clear how well it scales to high-dimensional problems, such as the Kolmogorov flow.
Our proposed method sits between stabilization techniques and learning invariant measures. In particular, we stabilize the training by implicitly learning the invariant measure of the system along the short-term dynamics.
MMD in Generative Modeling
@@
MMD-based regularization has been used in the context of generative modeling, e.g., the MMD has been used to distinguish between samples of the generated and true distributions (Li et al., 2015; Dziugaite et al., 2015; Li et al., 2017; Bińkowski et al., 2018) within the framework of generative adversarial network (Goodfellow et al., 2020). Additionally, drawing on the close connection between the MMD and proper scoring rules (Gneiting & Raftery, 2007; Ramdas et al., 2017), Si et al. (2021, 2023) use the energy distance (Baringhaus & Franz, 2004; Székely et al., 2004), a special case of the MMD (Sejdinovic et al., 2013), to train normalizing flow generative models (Rezende & Mohamed, 2015; Papamakarios et al., 2021).

### v2

@@ 4 Main Idea and Methods
To tackle the issue of distribution shift, we propose to focus on systems’ invariant measure preservation. Specifically, many systems of interest have some measure $\mu^{*}$ supported on an attractor that is invariant to the transformation ${\mathcal{S}}$ (Tucker, 2002; Weinan & Liu, 2002; Luzzatto et al., 2005; Ferrario, 2008; Hawkins, 2021). We cast our learning problem as finding parameters $\theta$ such that a surrogate ${\mathcal{S}}_{\theta}$ preserves $\mu^{*}$ while approximating ${\mathcal{S}}$ locally, which defines the following constrained optimization:
@@
where ${\mathcal{L}}(\theta)$ is the short-term loss, and $\mu_{\theta}^{*}$ is the invariant measure of ${\mathcal{S}}_{\theta}$, i.e., $({\mathcal{S}}_{\theta})_{\sharp}\mu_{\theta}^{*}=\mu_{\theta}^{*}$, which we assume exists [^7] This is a key hypothesis in our methodology. Otherwise, this assumption can be enforced in ${\mathcal{S}}_{\theta}$ by adding a potential term, as done in Li et al. (2022).. Solving this constrained optimization inherently alleviates the distribution shift problem. Since trivial solutions exist for measure preservation, e.g., if ${\mathcal{S}}_{\theta}$ is the identity, the trajectory matching component ${\mathcal{L}}(\theta)$ of this constrained objective is necessary for producing useful surrogates.
@@
Measure Distance   Our choice of measure distance needs to satisfy several desiderata, namely it should: i) respect the underlying geometry of ${\mathcal{U}}$ and support comparison between measures with non-overlapping supports, ii) admit an unbiased, sampled-based estimator, iii) have low computational complexity with respect to the system dimension and number of samples, iv) entail convergence properties on the space of measures defined on ${\mathcal{U}}$ (informally, $\textsc{D}(\mu_{\theta}^{*},\mu^{*})\rightarrow 0\implies\mu_{\theta}^{*}\rightarrow\mu^{*}$), and v) enjoy parametric rates of estimation (i.e., sampling error $|\textsc{D}-\widehat{\textsc{D}}|$ is independent of system dimension).
@@
Some popular notions of distance / divergence from statistical learning theory include the Kullback-Leibler and Hellinger. However, these do not take into account the distance metric of the space on which the distributions are defined (Genevay et al., 2018; Feydy et al., 2019) and in some cases are undefined for non-overlapping supports.
@@
In contrast, Integral Probability Metrics (IPMs; Müller (1997)) represent a general purpose tool for comparing two distributions. Among the class of IPMs, the Maximum Mean Discrepancy (MMD; Gretton et al. (2012)) stands out as it has a closed form expression and satisfies all our requirements described above. Deferring several details about the MMD to Appendix B, we define it here as,
$\mathrm{MMD}^{2}(\mu^{*},\mu_{\theta}^{*})=\mathbb{E}_{{\bm{u}},{\bm{u}}^{\prime}\sim\mu^{*}}[\kappa({\bm{u}},{\bm{u}}^{\prime})]$ (10)
$+\mathbb{E}_{{\bm{v}},{\bm{v}}^{\prime}\sim\mu_{\theta}^{*}}[\kappa({\bm{v}},{\bm{v}}^{\prime})]-2\mathbb{E}_{{\bm{u}}\sim\mu^{*},{\bm{v}}\sim\mu_{\theta}^{*}}[\kappa({\bm{u}},{\bm{v}})],$
@@
where $\kappa:{\mathcal{U}}\times{\mathcal{U}}\rightarrow\mathbb{R}$ is a kernel. [^8] The choice of kernel has important practical implications (Liu et al., 2020; Schrab et al., 2023), and many kernels $\kappa_{\sigma}$ are controlled by a bandwidth hyperparameter $\sigma$ that should be tuned. For two sets of $n$ samples $\{{\bm{u}}^{(i)}\}_{i=1}^{n}\sim\mu^{*}$ and $\{{\bm{v}}^{(i)}\}_{i=1}^{n}\sim\mu_{\theta}^{*}$, Equation 10 admits an unbiased estimator (Gretton et al., 2012):
$\widehat{\mathrm{MMD}}^{2}(\mu^{*},\mu_{\theta}^{*})=\frac{1}{n(n-1)}\Bigg{[}\sum_{i=1}^{n}\sum_{j\neq i}^{n}[\kappa({\bm{u}}^{(i)},{\bm{u}}^{(j)})]$
$+\sum_{i=1}^{n}\sum_{j\neq i}^{n}[\kappa({\bm{v}}^{(i)},{\bm{v}}^{(j)})]\Bigg{]}-\frac{2}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}[\kappa({\bm{u}}^{(i)},{\bm{v}}^{(j)})],$
@@
This estimator can be easily computed in ${\mathcal{O}}(n^{2})$ operations. Additionally, the MMD entails convergence properties on the space of measures defined on ${\mathcal{U}}$. That is, it metrizes weak convergence (Simon-Gabriel et al., 2023), and, for characteristic kernels, we have $\mathrm{MMD}(\mu^{*},\mu_{\theta}^{*})=0\iff\mu^{*}=\mu_{\theta}^{*}$ (Sriperumbudur et al., 2010). The MMD also enjoys parametric rates of estimation, with ${\mathcal{O}}(1/\sqrt{n})$ sampling error (Gretton et al., 2006; Tolstikhin et al., 2016). We point out that, in practice, $n$ is the batch size, since we employ stochastic optimization methods.
@@
Using the approximation in Equation 11 in the context of a stochastic optimization pipeline requires that we estimate this term with a potentially small batch size. In fact, besides some simple systems, one typically can only afford small batch sizes, which means we may not be fully capturing both $\mu^{*}$ and $\mu_{\theta}^{*}$. Although our choice of regularization loss comes with an unbiased estimator with parametric rates of error, in small regimes, estimation error can still diminish its effectiveness at providing a meaningful signal.
@@
$\mathrm{MMD}^{2}(({\mathcal{S}}^{k})_{\sharp}\mu^{*},({\mathcal{S}}_{\theta}^{k})_{\sharp}\mu^{*})=\mathbb{E}_{{\bm{u}},{\bm{v}}\sim\mu^{*}}[\kappa({\mathcal{S}}^{k}({\bm{u}}),{\mathcal{S}}^{k}({\bm{v}}))$
$\quad+\kappa({\mathcal{S}}_{\theta}^{k}({\bm{u}}),{\mathcal{S}}_{\theta}^{k}({\bm{v}}^{\prime}))-2\kappa({\mathcal{S}}^{k}({\bm{u}}),{\mathcal{S}}_{\theta}^{k}({\bm{v}})],$ (12)
which can be estimated by extracting a subset of initial conditions, time-evolving them $k$ steps, and then computing the estimator on the time-evolved samples. We point out that this is equivalent to conditioning the loss on the initial conditions, hence the name conditional regularization. When using samples to estimate MMD, we use a collection of initial conditions $\{{\bm{u}}_{0}^{(i)}\}_{i=1}^{n}$ for unconditional regularization and a collection of samples time-evolved by the true system, i.e., ones that come from later time steps in training trajectories, $\{{\bm{u}}_{k}^{(i)}={\mathcal{S}}^{k}({\bm{u}}_{0}^{(i)}))\}_{i=1}^{n}$ for conditional regularization.
@@
^$\widehat{{\mathcal{L}}}_{\lambda}^{\textsc{D}}(\theta)=\widehat{{\mathcal{L}}}^{\text{obj}}(\theta)+$ $\lambda_{1}\widehat{\textsc{D}}(\mu^{*},({\mathcal{S}}_{\theta}^{\ell})_{\sharp}\mu^{*})$ (13)
$+$ $\lambda_{2}\widehat{\textsc{D}}(({\mathcal{S}}^{\ell})_{\sharp}\mu^{*},({\mathcal{S}}_{\theta}^{\ell})_{\sharp}\mu^{*}),$
@@
For the measure distance $\widehat{\textsc{D}}$ in Equation 13, we use $\widehat{\mathrm{MMD}}^{2}$ and define $\kappa_{\bm{\sigma}}$ as a mixture of rational quadratic kernels (Rasmussen et al., 2006; Li et al., 2015):
@@ 5 Experiments
Evaluation   We evaluate models both for their ‘short-term’ predictive ability and ‘long-run’ stability. The former is measured by a cosine similarity statistic between true and predicted trajectories. The latter is measured by system-specific metrics (see Wan et al. (2023a)) capturing the distributional similarity between true and generated trajectories, along with their visual inspection. In particular, we use the Sinkhorn Divergence (SD) (Genevay et al., 2018) between the empirical distributions of ground truth and predicted trajectories at various time steps to quantify distributional overlap. We also use the mean energy log ratio (MELR), which measures the average deviation of the energy at each Fourier mode of the generated snapshots when compared to the ground truth. We also consider its weighted variant (MELRw), which up-weights modes with higher energy, in particular the low-frequency modes. We also use the mean of the Frobenius norm of covariance matrix (covRMSE) that measures the spatial statistical properties of the generated samples. We additionally compute point-wise Wasserstein metrics. Finally, we consider a time correlation metric (TCM) which provides a measure of temporal behavior, in contrast to most of the metrics above, which are snapshot-based. For more detail on and precise definitions of the evaluation criteria, see Appendix C.
@@ 5.1 Lorenz 63
Figure 1 (b) and (c) demonstrate the improved stability from adding invariant measure regularization to the different training objectives considered in this paper. In particular, Figure 1 (b) demonstrates improved long-term statistics, as the distribution of points for models trained with DySLIM are closer to that of ground truth trajectories compared to those from models trained with unregularized objectives. In addition, Figure 1 (c) shows that with DySLIM we obtain the added benefits of improving the short-term model prediction, with longer de-correlation times. For further results with different metrics, see Appendix F.1.
@@ 5.2 Kuramoto-Sivashinsky
In Figure 2, we observe better short-term predictions and improved long-term stability, as measured by lower SD between ground truth and predicted trajectory distributions. In Figure 1 (a), we see example trajectories that highlight the difference between models trained with and without regularized objectives. For the Curr objective, models often diverge and produce numerical instability, while for the Pfwd objective, models deviate from the attractor. In contrast, the regularized versions of these objectives yield more stable models that remain on the correct attractor manifold.
@@ 5.3 Kolmogorov Flow
Finally, we study chaotic 2D fluid flow defined by the Navier-Stokes equations with Kolmogorov forcing. Information about the PDE and experiment setups is available in Appendix E.3. For this system, the SD becomes non-discriminative due to the high-dimension of the state space, so we rely on the other metrics outlined above.
Figure 3 (left) shows typical behavior of the unrolled trained models for a given initial condition: the baselines become highly dissipative and quickly veer towards the mean, whereas DySLIM greatly improves long-term behavior. Figure 3 (right) shows a similar results to those in the other experiments: the short-term behavior of the solution is enhanced by the regularization (see Figure 8 for further comparisons). We find that curriculum training is often worse due to more stringent memory requirements that prevent us from unrolling for longer time-horizons during training.
As an ablation, we sweep over different batch sizes and learning rates (see Appendix E.3 for the specific details.) The results are summarized in Table 1, which shows that models trained with DySLIM either have an edge or remain competitive across the spectrum of different learning rates and batch sizes considered. However, as batch size and learning rate increase, the behavior of the model trained with DySLM remains consistent, whereas the models trained only using the original objective deteriorate quickly.
@@
Figure 3: (Left) Sample reference and predicted trajectory across models trained on the Kolmogorov Flow data using the Curr and Pfwd objectives, together with the regularized versions. (Right) Evolution of the cosine similarity over time for Curr and Pfwd objectives with and without regularization. The solid line is the median among $160$ runs ($32$ trajectories for each of the $5$ random seeds), and the shaded regions correspond to the second and third quartile. ($\lambda_{1}=0$, $\lambda_{2}=100$, batch size = 128 and learning rate = $5\mathrm{e}^{-4})$.
Table 1: Kolmogorov flow: Metrics for 1-step, curriculum, and pushforward objectives without and with regularization ($\lambda_{1}=0$, $\lambda_{2}=100$). Boldface numbers indicate that the metric is improved by our regularization. All values displayed are in units of ×$\times 10^{-2}$.
Batch size LR
MELR ($\downarrow$)
MELRw ($\downarrow$)
covRMSE ($\downarrow$)
Wass1 ($\downarrow$)
TCM ($\downarrow$)
@@ 6 Related Work
ROM Methods   Classical reduced order model (ROM) methods build surrogates by identifying low-dimensional linear approximation spaces tailored to representing target system states. Such spaces are usually derived from data samples (Aubry et al., 1988; Barrault et al., 2004; Chinesta et al., 2011; Amsallem et al., 2012), and ROMs are obtained by projecting the system equations onto the approximation space (Galerkin, 1915). Although these methods inherently leverage the linear behavior of underlying dynamics, they have been recently extended to handle mildly non-linear dynamics (Willcox, 2006; Astrid et al., 2008; Chaturantabut & Sorensen, 2010; Ayed et al., 2019; Geelen et al., 2022). However, their performance deteriorates rapidly for highly non-linear advection-dominated systems, such as KS and Kolmogorov flow (Peherstorfer, 2022)
@@
Hybrid Physics-ML   More recent methods hybridize classical numerical methods with contemporary data-driven deep learning techniques (Mishra, 2018; Bar-Sinai et al., 2019; Bruno et al., 2021; Kochkov et al., 2021; List et al., 2022; Frezat et al., 2022; Dresdner et al., 2022; Boral et al., 2023). These approaches learn corrections to numerical schemes from high-resolution simulation data, resulting in fast, low-resolution methods with high-accuracy. However, they require knowledge of the underlying PDE.
@@
Pure ML-surrogates and Stabilization Techniques   Recent works have focused on short-term training trajectories using recurrent networks (Vlachas et al., 2018) and reservoir computing techniques (Vlachas et al., 2020; Platt et al., 2021). Other approaches seek to regularize the training stage by leveraging properties of the systems. Such stabilization can be achieved by incorporating noise (Sanchez-Gonzalez et al., 2020), which can be induced by the learned model (Brandstetter et al., 2022); by back-propagating the gradient along many time steps (Um et al., 2020), and learning the dynamics on a latent space (Stachenfeld et al., 2022; Serrano et al., 2023), while promoting smoothness in the latent space (Wan et al., 2023b). Or by using a generative teacher network (Lamb et al., 2016), or leveraging an approximate inertial form (Lu et al., 2017). Related to the Curriculum training baseline, Hess et al. (2023) use states that interpolate between model predicted and ground truth states to mitigate gradient explosion. Finally, a somewhat related stabilization method, introduced in Wang et al. (2014); Blonigan et al. (2018), develops a shadowing technique for sensitivity analysis of long-term averaged gradients.
@@
Operator Learning   Neural operators seek to learn the integro-differential operators directly, without explicit PDE-informed components.These methods often leverage classical fast-methods (Fan et al., 2019; Li et al., 2020, 2021; Tran et al., 2021), or approximation-theoretic structures (Lu et al., 2021) to achieve computational efficiency. Some of these techniques have been extended to handle dissipative systems (Li et al., 2022) by hard-coding a dissipative term at both training and inference time.
Table 2: Complexities for each objective, with $d$ denoting state dimension, $|\theta|$ number of parameters, which is implementation dependent, $NN$ complexity of the neural network for one application, $n_{b}$ batch size, and $n_{t}$ number of maximum rollout steps.
Objective Cost $\mathcal{O}(\cdot)$ Memory footprint $\mathcal{O}(\cdot)$
1-step $n_{b}d+n_{b}NN$ $n_{b}d+n_{b}|\theta|$
+ DySLIM $n_{b}^{2}d+n_{b}NN$ $n_{b}^{2}d+n_{b}|\theta|$
Curr $n_{t}n_{b}d+n_{t}n_{b}NN$ $n_{t}n_{b}d+n_{t}n_{b}|\theta|$
+ DySLIM $n_{t}n_{b}^{2}d+n_{b}NN$ $(n_{b}^{2}+n_{b})n_{t}d+n_{t}n_{b}|\theta|$
Pfwd $n_{b}d+n_{b}NN$ $n_{b}d+n_{b}|\theta|$
+ DySLIM $n_{b}^{2}d+n_{b}NN$ $n_{b}^{2}d+n_{b}|\theta|$
@@
Learning Invariant MeasuresBotvinick-Greenhouse et al. (2023) use an Eulerian approach to learn dynamics on invariant measures for low-dimension ODEs using the Feynman-Kac formula coupled with PDE-constrained optimization. In recent work, Jiang et al. (2023) use neural operators and optimal transport to match the distribution of system-specific summary statistics, which are built using knowledge of the underlying equation driving the system dynamics. However, the approach was only applied to small systems with low-dimensional attractors, and it is not clear how well it scales to high-dimensional problems, such as the Kolmogorov flow. Similarly, Platt et al. (2023) regularize using other invariants, such as the Lyapunov spectrum.
Our proposed method sits between stabilization techniques and learning invariant measures. In particular, we stabilize the training by implicitly learning the invariant measure of the system along the short-term dynamics. While our work elects to use the MMD, we note that using other Optimal Transport-based metrics for measure matching, such as the Sinkhorn Divergence, as in (Jiang et al., 2023), is a reasonable choice when the system state is small, and batch size is large, as shown in Appendix G. However, as system size increases and batch sizes decrease (due to memory constraints), we find that the models trained using SD in the regularization perform worse compared to those trained using MMD.
@@
MMD in Generative Modeling   MMD-based regularization has been used in the context of generative modeling, e.g., the MMD has been used to distinguish between samples of the generated and true distributions (Li et al., 2015; Dziugaite et al., 2015; Li et al., 2017; Bińkowski et al., 2018) within the framework of generative adversarial network (Goodfellow et al., 2020). Additionally, drawing on the close connection between the MMD and a related proper scoring rule (Gneiting & Raftery, 2007; Ramdas et al., 2017), Si et al. (2021, 2023) use the energy distance (Baringhaus & Franz, 2004; Székely et al., 2004), a special case of the MMD (Sejdinovic et al., 2013), to train normalizing flow generative models (Rezende & Mohamed, 2015; Papamakarios et al., 2021).
Complexity   Our methodology incurs relatively small overhead compared to the baselines. Table 2 shows that our methodology adds an extra cost depending only quadratically on the batch size and linearly on state dimension. We see this empirically in Table 3: wall-clock times are roughly equal with and without our regularization.
Table 3: Median execution times (rounded to the hour) for 5 training runs (720k steps) on Kolmogorov flow using the PushFwd (max rollout of 10) and Curr (max rollout of 5).
Batch size
Pfwd
Curr
Base DYSLIM Base DYSLIM
32 40 40 48 48
64 74 76 161 163
128 145 146 OOM OOM